ID |
Date |
Author |
Subject |
Text |
 |
344
|
Tue Mar 1 08:02:26 2022 |
SHB, MS | check |
'Check Clock Status', 'Check
ADC Calibration' and 'Check the White |
11x |
343
|
Tue Mar 1 06:35:42 2022 |
SHB, MS | check | detector motors are still out
'Check Clock Status', 'Check
ADC Calibration' : Clock status test |
11x |
342
|
Tue Mar 1 06:04:04 2022 |
SHB, MS | positions | just FYI: current positions saved in files
are x=1000 and y=-750. this information is
good in case of crash. |
|
341
|
Tue Mar 1 05:23:02 2022 |
JM, EM | Current status | 06:20
Servo motors moved to fully out
position as some work on the beam was neccessary. |
|
340
|
Tue Mar 1 04:07:31 2022 |
JM, EM | Temperature cycle | Have been unable to achieve decent target
density with temperature cycling so far.
Will attempt to go up to 300 k and back down |
|
339
|
Tue Mar 1 03:34:28 2022 |
JM, EM | Current status | Temperature setpoint = 116 k
Detector IN potentiometer = 34.89
mm |
7x |
338
|
Tue Mar 1 03:19:50 2022 |
JM, EM | Temperature Cycle | 04:20
Did another temperature cycle (118
K -> 125 K -> 95 K -> 116 K) trying |
|
337
|
Tue Mar 1 02:15:29 2022 |
JM, EM | Current status | 03:15
The nozzle motor control at the
target crashed, but operation has been restored. |
6x |
336
|
Mon Feb 28 23:54:02 2022 |
JM, EM | Current Status, Temperature cycle | 00:51
Temperature setpoint = 116 K
Detector position IN potentionmeter |
  |
335
|
Mon Feb 28 22:34:51 2022 |
JM, EM | Current status | 23:35
Temperature setpoint = 116 K
Detector position IN, potentiometer |
    |
334
|
Mon Feb 28 14:47:52 2022 |
CB, TD | Monday 28 February 15:00-23:00 |
15.51 1.8.W spectra 20ms, 2ms, 200us, 20us
FSR - attachments 1-8
|
17x |
333
|
Mon Feb 28 14:42:48 2022 |
CB, TD | Cryostat set to 115K |
Cryostat set to 115K |
|
332
|
Mon Feb 28 14:38:15 2022 |
TD | NewMerger statistics |
Screenshots of all NewMerger statistics |
37x |
331
|
Mon Feb 28 13:56:02 2022 |
JG, CB | Temperature cycle to 125K | 14:45
Rutherford rate was decreased significantly
for longer period. |
|
330
|
Mon Feb 28 13:16:39 2022 |
JG, AK, CB | Current status | 14:16
beam intensity at injection: 7e7
ions |
8x |
329
|
Mon Feb 28 11:57:22 2022 |
JG, AK, CB | Current status | 12:55
beam intensity at injection: 9e7
ions |
8x |
328
|
Mon Feb 28 10:55:48 2022 |
JG, US | Current status | 11:55
beam intensity at injection: 8e7
ions |
8x |
327
|
Mon Feb 28 09:31:48 2022 |
JG, AK, CB | Current status | 10:35
beam intensity at injection: 1e8
ions |
8x |
326
|
Mon Feb 28 08:36:35 2022 |
JG, AK, CB | Current status | 09:35
beam intensity at injection: 8.2e7
ions |
8x |
325
|
Mon Feb 28 07:35:51 2022 |
JG | current logger crashed | The current logger did crash in the evening
of 25.02. and stopped logging since then,
it seems. |
|
|